A solution is made by combining 15.0 mL of 16.5 M acetic acid with 5.50 g of sodium acetate and diluting to a total volume of 1.50 L. What is the pH of the solution?
1
Expert's answer
2020-05-13T14:20:31-0400

C=n/v=>n(CH3COOH)=c*v=16,5*0, 015=0,2475 moll

n(CH3COONa)=m/Mr=5,5/82=0,067 moll

c2(CH3COOH)=n/v=0,2475/1,5=0,165

c(CH3COONa)=n/v=0,067/1,5=0,045

pH=pK(CH3COOH)+lg(c(CH3COONA)/c(CH3COOH))=4,75+lg(0,045/0,165)=4,19